OpenStudy (anonymous):

What is the value of x in the equation x – y = 30, when y = 15?

OpenStudy (anonymous):

We have x - 15 = 30

OpenStudy (anonymous):

Add 15 + 30

OpenStudy (anonymous):

x - y = 30 Plug in 15: x - 15 = 30 Add 15 to both sides, what's 30 + 15?

OpenStudy (anonymous):

substitute 15 in for y.

OpenStudy (anonymous):

45 LOL

OpenStudy (anonymous):

Yes, that's our answer.

OpenStudy (anonymous):

45 is not the answer lol

OpenStudy (anonymous):

I think im missing something

OpenStudy (anonymous):

Oh, your equation must be wrong..